40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 24 22:55:50 2019, gautam, Update, Electronics, Satellite box S/N 102 investigation IMG_7294.JPGD961289-B2.pdf
    Reply  Tue Jan 29 17:19:16 2019, gautam, Update, Electronics, Satellite box S/N 105 repaired IMG_7301.JPG
Message ID: 14421     Entry time: Tue Jan 29 17:19:16 2019     In reply to: 14417
Author: gautam 
Type: Update 
Category: Electronics 
Subject: Satellite box S/N 105 repaired 

[chub, koji, gautam]

Attachment #1 shows the signal routing near the Satellite box. Somehow, the female 64 pin IDC connector that brings the signals from the coil driver board wasn't mating well with the mail connector on the Satellite box front panel. This is a connector specific problem - plugging the female end into one of the male connectors inside the Satellite box yielded signal continuity. The problem was resolved by re-making both connections -by driving the EPICS bias slider through its full range, we were able to see the full voltage swing at the DB connectors going to the flange

This kind of flakiness could be all around the lab, and could be responsible for many of the suspension "mysteries". To re-iterate, the problem seems to be the way the female sockets of the connector mates with the male pins - while the actual crimping points may look secure, there may not be signal continuity.

Now that this problem is resolved, tomorrow we will recover the cavity alignment and possibly start a pumpdown.


Unrelated to this work - the spare satellite box (S/N #100), which had a note on it that said "low voltages", was tested. The "low voltages" referred to the OSEM shadow sensor voltages being low when the LED was completely unobscured. The reason was that the mod to increase the drive current to 25 mA had not yet been implemented on this unit. I added the appropriate 806 ohm resistors, and verified that the voltages were correct, so now we have a working spare. It is stored in the "photodiode" cabinet along the east arm, together with the tester boxes. 

Attachment 1: IMG_7301.JPG  608 kB  | Hide | Hide all
IMG_7301.JPG
ELOG V3.1.3-